OpenStudy (anonymous):

Shania is having trouble understanding her algebra assignment. To help her understand, she asks Mark why is not equal to . Which of the following responses would be best for Mark to give? A. “You should multiply the exponents rather than add them. The correct answer is .” B. “The x and y are in the wrong order. The correct answer is .” C. “The multiplication property of exponents only holds when the numbers being multiplied have the same base. Since the bases are different, the expression cannot be simplified any further.” D. “

OpenStudy (yanasidlinskiy):

looks like C to me can't mix x's and y's x^2*x^3 = x*x * x*x*x = x^5 but x^2 * y^3 = x*x * y*y*y